# Riemannian Geometry of Optimal Rebalancing in Dynamic Weight Automated Market Makers

### 09 Mar 2026 v2

@@ 7 Experiments
We compare interpolation methods numerically using the same $N=3$ token setup as [1]: $\boldsymbol{\mathbf{w}}(t_{0})=\{0.05,0.55,0.4\}$, $\boldsymbol{\mathbf{w}}(t_{f})=\{0.4,0.5,0.1\}$, $f=1000$ steps.
@@
Weight trajectories and block-to-block changes are plotted in Appendix G (Figures G.2 and G.2). At this scale, (AM+GM)/normalise and SLERP are visually indistinguishable, consistent with the exact midpoint equivalence (Theorem 2) and third-order agreement at general tt5.1). For SLERP, block-to-block changes vary smoothly; constant metric speed in the Fisher–Rao metric does not correspond to constant weight increments in Cartesian coordinates, but the variation is small ($\sim 10^{-4}$).
@@
Figure 5: Per-step arbitrage loss $\log r_{k}$ across 1000 steps, $N=3$ setup. Left to right: linear (std/mean =0.32=0.32), (AM+GM)/normalise ($0.086$), SLERP ($0.0002$). SLERP achieves near-perfect uniformity, orders of magnitude better than other methods.
@@
Table 2: Per-step loss uniformity: ratio of standard deviation to mean of $\log r_{k}$.
@@
SLERP achieves near-perfect loss uniformity, Table 2, consistent with it traversing the Fisher–Rao metric geodesic at constant speed. The residual variation of $0.02\%$ arises from higher-order terms beyond the quadratic approximation of Eq (10).
@@
To verify that SLERP is near-optimal on the exact KL cost, we ran L-BFGS-B optimisation over all f1f-1 intermediate weight vectors, initialised from the SLERP solution, for the $N=3$ setup. At $f=1000$, the optimiser converged in 29 iterations without improving the objective beyond SLERP’s value. At $f=50$, the optimiser achieved a retained fraction of 0.989 048 060.989\,048\,06 versus SLERP’s 0.989 047 930.989\,047\,93, a relative cost improvement of ${\sim}1.2\times 10^{-5}$, consistent with the $O(\Omega^{2}/f^{2})$ bound of Theorem 3.
@@
Figure 6: Weight differences between interpolation methods. Left: SLERP minus linear ($\sim 0.03$), reflecting the non-linearity of the geodesic. Centre: SLERP minus (AM+GM)/normalise ($\sim 0.003$), an order of magnitude smaller, with an S-curve vanishing at both endpoints and the midpoint, consistent with Theorem 2. Right: brute-force optimal minus SLERP ($\sim 10^{-7}$), confirming SLERP is near-optimal.
Figure 6 confirms that the SLERP-minus-linear differences are large ($\sim 0.03$) with the same arch shape as the optimal-minus-linear plot in [1]. The SLERP-minus-(AM+GM)/normalise differences are an order of magnitude smaller ($\sim 0.003$), and the differences in weights between SLERP and the brute force optimal trajectory are of order $10^{-7}$.
Convergence across step counts
Figure 7: $N=3$ near-boundary configuration $\boldsymbol{\mathbf{w}}:(0.01,0.01,0.98)\to(0.49,0.49,0.02)$. Left: total arbitrage loss vs. step count $f$ for each interpolation method and the brute-force L-BFGS-B optimum. Right: relative sub-optimality (ratio of method loss to brute-force optimal loss minus one). SLERP closely tracks the brute-force optimum across all $f$; linear interpolation diverges by up to ${\sim}20\%$ near the boundary.
Figure 7 shows total loss and relative sub-optimality as a function of the step count $f$ for the $N=3$ near-boundary configuration. SLERP closely tracks the brute-force optimal across steps, validating the $O(\Omega^{2}/f^{2})$ sub-optimality bound of Theorem 3. Linear interpolation diverges from the optimum as $f$ grows, incurring up to ${\sim}20\%$ excess loss at this near-boundary configuration. (AM+GM)/normalise performs slightly worse than SLERP but substantially better than linear, consistent with the third-order agreement established in §5.1.
@@
Current G3M implementations enforce a minimum weight of $1\%$, so the practical boundary is $w_{\min}=0.01$. At this floor the Fisher–Rao metric diverges and the quadratic approximation underlying SLERP is least accurate, making near-boundary performance an important stress test.
Figure 8: Left: total loss at $f=1000$ as a function of the smallest weight $w_{\min}$, for $N=2$ symmetric weight changes $(w_{\min},\,1{-}w_{\min})\to(1{-}w_{\min},\,w_{\min})$. Right: loss ratio relative to SLERP. SLERP’s advantage over both linear and (AM+GM)/normalise grows as weights approach the boundary.
Figure 8 shows that SLERP’s advantage over both linear and (AM+GM)/normalise widens as weights approach the boundary: at $w_{\min}=0.01$, linear interpolation incurs ${\sim}20\%$ more loss than SLERP, and (AM+GM)/normalise incurs ${\sim}3.5\%$ more. For interior weights ($w_{\min}\geq 0.2$), all three methods are within $1\%$ of each other.
Figure 9: Per-step loss $\log r_{k}$ near the simplex boundary ($N=3$, $f=1000$, $\boldsymbol{\mathbf{w}}:(0.01,0.01,0.98)\to(0.49,0.49,0.02)$). SLERP achieves near-constant loss per step (std/mean =0.0011=0.0011); other methods show large variation, especially at steps where weights pass through their smallest values.
Figure 9 shows per-step losses for the $N=3$ near-boundary configuration at $f=1000$. SLERP maintains near-perfect uniformity (std/mean =0.0011=0.0011) even with weights at the $1\%$ floor: compare with std/mean =0.0002=0.0002 for the interior configuration of Figure 5. Linear interpolation shows extreme non-uniformity (std/mean =0.89=0.89) with a loss spike at the boundary-crossing steps. Appendix H provides further near-boundary analysis, including a comparison of single-midpoint ($f=2$) methods where the exact Lambert W formula of [1] outperforms SLERP by up to $10\%$.
8 Discussion
@@
In the language of information geometry, the arbitrage cost is the canonical divergence of the dually-flat simplex (Appendix E). The Fisher–Rao metric provides a single framework that explains why the various candidate interpolation methods (linear, geometric, (AM+GM)/normalise, Lambert W) all perform similarly. Each corresponds to a different geodesic or approximate geodesic on the weight simplex (Appendix E), and their near-agreement reflects the fact that the simplex geometry is nearly flat at the scale of typical per-step weight changes. The sub-optimality bound (Theorem 3) makes this precise: SLERP’s gap to the exact KL optimum is $O(\Omega^{4}/f^{3})$, so for any reasonable step count the choice between methods is practical, not a matter of loss optimality.
Constant-price assumption
The analysis assumes market prices are constant within each interpolation (§2). In reality prices change over time, and arbitrage may be partial or delayed, introducing additional effects outside the KL framework. The optimal interpolation under time-varying prices remains an open question.
Multi-token generalisation
The Lambert W approach of [1] optimises each weight component independently and renormalises, which is only exact for small weight changes. SLERP operates intrinsically on the sphere and automatically preserves the constraint $\sum_{i}w_{i}=1$ (since $\sum_{i}\eta_{i}^{2}=1$ along any great circle in the positive orthant), making it a natural NN-token generalisation with no renormalisation required.
Fees and transaction costs
The analysis assumes zero fees and perfect arbitrage at every block. In practice, fees and gas costs impose a profitability threshold: arbitrage occurs only when the accumulated opportunity exceeds transaction costs, reducing the effective rebalancing frequency from once per block to once every several minutes or hours [8]. At this coarser cadence the number of effective steps $f$ is small, so the absolute cost of rebalancing is dominated by the unavoidable $2\Omega^{2}/f$ term (Theorem 3) rather than by the choice of interpolation method.
@@
Computing SLERP requires one $\arccos$ evaluation, NN square roots, and per-step $\sin$ calls, comparable in cost to geometric mean interpolation. For on-chain implementations where transcendental functions are expensive, the (AM+GM)/normalise heuristic of [1] remains the most practical choice. The recursive bisection algorithm (Corollary 3) shows that this heuristic computes exact geodesic points at power-of-two step counts using only elementary arithmetic. No trigonometric functions are needed.
Accuracy of the quadratic approximation
The quadratic loss kernel, Eq (10), has cubic remainder in the per-step weight changes: for $f$-step interpolation the cumulative error is $O(1/f^{2})$. For large single-step updates or weights near zero, the Lambert W approach of [1] (which maximises the exact finite-step retention ratio) may be preferred.
9 Concluding remarks
The per-step arbitrage loss from rebalancing a TFMM pool is a KL divergence; the Fisher–Rao metric therefore governs rebalancing costs on the weight simplex. The loss-minimising interpolation under the leading-order expansion of this cost is a geodesic traversed at constant speed (SLERP in Hellinger coordinates), and the (AM+GM)/normalise heuristic of [1] computes exact geodesic midpoints via a binomial identity. The recursive bisection algorithm (Corollary 3) produces the full SLERP trajectory at power-of-two step counts using only elementary arithmetic. The sub-optimality bound (Theorem 3) guarantees near-optimality on the exact KL cost. The (AM+GM)/normalise heuristic remains the practical on-chain recommendation.
The geometric structure here is not specific to G3M/TFMM pools. Any AMM whose trading function is parameterised by a vector on the probability simplex will have its arbitrage cost naturally expressible as a divergence in the information-geometric sense, making the Fisher–Rao metric the natural Riemannian structure. Different AMM mechanisms (e.g. virtual balance interpolation in concentrated liquidity designs) would induce different, potentially state-dependent, metrics on their parameter spaces. Identifying these metrics and their geodesics could yield analogous rebalancing strategies for other designs.
@@ Appendix A Retention Ratio
A.1 Deferred Proof: Proposition 1 [Retention ratio]
@@
$R_{i}^{\mathrm{end}}=R_{i}^{\mathrm{start}}\frac{w_{i}^{\mathrm{end}}}{w_{i}^{\mathrm{start}}}\prod_{j=1}^{N}\left(\frac{w_{j}^{\mathrm{start}}}{w_{j}^{\mathrm{end}}}\right)^{w_{j}^{\mathrm{end}}},$ (A.1)
@@
$\frac{R_{i}^{\mathrm{end}}}{R_{i}^{\mathrm{start}}}\frac{w_{i}^{\mathrm{start}}}{w_{i}^{\mathrm{end}}}=\prod_{j=1}^{N}\left(\frac{w_{j}^{\mathrm{start}}}{w_{j}^{\mathrm{end}}}\right)^{w_{j}^{\mathrm{end}}},$ (A.2)
and we thus define $r:=\prod_{j=1}^{N}\left(\frac{w_{j}^{\mathrm{start}}}{w_{j}^{\mathrm{end}}}\right)^{w_{j}^{\mathrm{end}}}$. For given reserves and market prices $\boldsymbol{\mathbf{m}}_{p}$, pool value is $V=\sum_{i=1}^{N}R_{i}m_{p,i}$. By the action of arbitrageurs, the pool holds the minimum value possible under its trading function. The weight vector then gives the division of value between assets [2], so we have for all $i$
$V^{\mathrm{start}}w^{\mathrm{start}}_{i}$ $=R^{\mathrm{start}}_{i}m_{p,i}$ (A.3)
$V^{\mathrm{end}}w^{\mathrm{end}}_{i}$ $=R^{\mathrm{end}}_{i}m_{p,i}$ (A.4)
@@
$\frac{V^{\mathrm{end}}}{V^{\mathrm{start}}}\frac{w^{\mathrm{end}}_{i}}{w^{\mathrm{start}}_{i}}$ $=\frac{R^{\mathrm{end}}_{i}}{R^{\mathrm{start}}_{i}}$ (A.5)
$\Rightarrow\frac{V^{\mathrm{end}}}{V^{\mathrm{start}}}$ $=\frac{R_{i}^{\mathrm{end}}}{R_{i}^{\mathrm{start}}}\frac{w_{i}^{\mathrm{start}}}{w_{i}^{\mathrm{end}}}$ (A.6)
$\Rightarrow\frac{V^{\mathrm{end}}}{V^{\mathrm{start}}}$ $=\prod_{j=1}^{N}\left(\frac{w_{j}^{\mathrm{start}}}{w_{j}^{\mathrm{end}}}\right)^{w_{j}^{\mathrm{end}}}$ (A.7)
$\Rightarrow\frac{V^{\mathrm{end}}}{V^{\mathrm{start}}}$ $=r$ (A.8)
@@
A.2 Deferred Proof: Corollary 1 [Quadratic loss kernel]
See 1

### 31 Mar 2026 v3

@@ 7.1. Constant prices
We compare interpolation methods numerically using the same $N=3$ token setup as [Willetts and Harrington, 2024a]: $\boldsymbol{\mathbf{w}}(t_{0})=\{0.05,0.55,0.4\}$, $\boldsymbol{\mathbf{w}}(t_{f})=\{0.4,0.5,0.1\}$, $f=1000$ steps.
@@
Weight trajectories and block-to-block changes are plotted in Appendix H (Figures H.3 and H.3). At this scale, (AM+GM)/normalise and SLERP are visually indistinguishable, as expected from the exact midpoint equivalence (Theorem 6) and third-order agreement at general tt3.7). For SLERP, block-to-block changes vary smoothly; constant metric speed in the Fisher–Rao metric does not correspond to constant weight increments in Cartesian coordinates, but the variation in increment size is small ($\sim 10^{-4}$ relative to the mean increment).
@@
Figure 6. Per-step arbitrage loss $\log r_{k}$ across 1000 steps, $N=3$ setup. Left to right: linear (std/mean =0.32=0.32), (AM+GM)/normalise ($0.086$), SLERP ($0.0002$). SLERP achieves near-perfect uniformity, orders of magnitude better than other methods.
@@
Table 2. Per-step loss uniformity: ratio of standard deviation to mean of $\log r_{k}$.
@@
SLERP achieves near-perfect loss uniformity (Table 2): constant-speed traversal of the geodesic equalises the per-step cost by construction. The residual $0.02\%$ variation comes from higher-order terms beyond the quadratic approximation (Eq (10)).
(a) Two-token pendulum (Corollary 2), $(0.5,0.5)\to(0.7,0.3)$, varying $\mu=fT\sigma^{2}/16$. Left: $\theta$-space trajectories. Centre: weight-space trajectories. Right: per-step KL and LVR cost decomposition for $\mu=0.004$ and $\mu=1$. At practical parameters ($\mu\ll 1$), the trajectory is indistinguishable from SLERP.
(b) Left: expected log-loss vs. step count at $\sigma=50\%$ ann. (10 00010\,000 GBM paths per point); the analytical $C(f)$ (Eq (25)) matches the MC estimate, and the constant-price curve $2\Omega^{2}/f$ underestimates above $f^{*}$. Right: $f^{*}$ vs. volatility; predicted values (dotted vertical lines) match the simulated optima, confirming $f^{*}\propto 1/\sigma$.
(c) Combined-objective (KL + LVR) optimisation, $\sigma=50\%$ ann. Left: SLERP and L-BFGS-B trajectories at $f=2000$ ($\lambda=9.0$). Centre: optimised\,-\,SLERP deviation (max $|\Delta w|=1.16\times 10^{-3}$). Right: relative improvement of L-BFGS-B over SLERP at three $(\!f,\lambda)$ operating points.
Figure 7. Stochastic-price experiments, two-token setup $(0.5,0.5)\to(0.7,0.3)$, $\sigma=50\%$ ann.
Figure 8. Historical backtest: ETH/USD, $(0.5,0.5)\to(0.7,0.3)$, across three market regimes (calm, volatile, trending). Dashed curves show the analytical $C(f)$; dots show realised $\log(V_{T}/V_{0})$ with 95% bootstrap confidence intervals. Vertical lines mark $f^{*}$ (red, dashed) and the constant-price $2\Omega^{2}/f$ reference. The GBM prediction matches realised costs; $f^{*}$ sits near the observed minimum in each regime.
@@
To verify that SLERP is near-optimal on the exact KL cost, we ran L-BFGS-B [Byrd et al., 1995] optimisation over all f1f-1 intermediate weight vectors, initialised from the SLERP solution, for the $N=3$ setup. At $f=1000$, the optimiser converged in 29 iterations without improving the objective beyond SLERP’s value. At $f=50$, the retained fraction improved from SLERP’s 0.989 047 930.989\,047\,93 to 0.989 048 060.989\,048\,06, a relative gain of ${\sim}1.2\times 10^{-5}$ (compare the $O(\Omega^{2}/f^{2})$ bound of Theorem 9).
@@
SLERP-minus-linear weight differences are $\sim 0.03$, with the same arch shape as the optimal-minus-linear plot in [Willetts and Harrington, 2024a]. SLERP-minus-(AM+GM)/normalise differences are an order of magnitude smaller ($\sim 0.003$), and differences between SLERP and the brute-force optimal trajectory are of order $10^{-7}$ (Figure H.1, Appendix H).
@@
Current G3M implementations enforce a minimum weight of $1\%$, where the Fisher–Rao metric diverges and the quadratic approximation is least accurate. SLERP’s advantage over other methods in fact widens near the boundary: at $w_{\min}=0.01$, linear interpolation incurs ${\sim}20\%$ more loss, and (AM+GM)/normalise ${\sim}3.5\%$ more. For interior weights ($w_{\min}\geq 0.2$), all three methods are within $1\%$ of each other. Full near-boundary experiments (convergence across step counts, per-step uniformity, and single-midpoint comparisons where Lambert W outperforms SLERP by up to $10\%$) are in Appendix I. Near the boundary, varying speed along the SLERP curve (fast through high-LVR regions, slow near vertices) reduces the total rebalancing-plus-LVR cost $C(f^{*})$ by ${\sim}3\%$; see Appendix L.
7.2. Stochastic prices
We use the two-token setup $(0.5,0.5)\to(0.7,0.3)$ with $\Delta t_{\mathrm{block}}=12$ s and annualised volatility $\sigma=50\%$ unless stated otherwise.
Pendulum trajectories
Figure 7(a) shows solutions of the two-token pendulum (Corollary 2) for $\mu$ from 0 (free geodesic) to $100$ (boundary-layer regime). At $\mu=0$ the trajectory is SLERP; as $\mu$ increases, the LVR potential deflects the path away from equal weighting toward the vertices (cf. Appendix K). At $\mu=100$ the trajectory jumps to a vertex and remains there for the bulk of the interpolation. At $\mu=0.004$ (the practical regime at $f=f^{*}$), the KL cost dominates and the LVR contribution is nearly flat; at $\mu=1$ the LVR cost varies significantly along the path.
Monte Carlo validation of $f^{*}$
Figure 7(b) compares the analytical cost $C(f)$ against Monte Carlo estimates (10 00010\,000 GBM paths per point). The predicted $f^{*}_{E[V]}=2744$ coincides with the observed minimum. At $\sigma=30\%$, $50\%$, and $80\%$ annualised, the predicted $f^{*}$ values (4574, 2744, 1715) agree with the simulated optima, as the $f^{*}\propto 1/\sigma$ scaling of Proposition 3 requires.
SLERP near-optimality under KL + LVR
Figure 7(c) compares SLERP against L-BFGS-B [Byrd et al., 1995] optimisation of the combined KL + LVR objective. At $f=2000$ ($\lambda=9.0$), the optimised trajectory deviates from SLERP by at most $|\Delta w|=1.16\times 10^{-3}$, the order of magnitude predicted by the Green’s function correction (Appendix K). The relative cost improvement over SLERP is ${\sim}10^{-7}$ at $f=50$ ($\lambda=0.01$), ${\sim}6\times 10^{-7}$ at $f=500$ ($\lambda=0.56$), and $1.2\times 10^{-4}$ at $f=2000$ ($\lambda=9.0$). At $f=f^{*}$, where $\lambda^{*}=O(1)$, there is nothing to gain from departing from SLERP.
Historical backtest
Figure 8 replays the rebalancing on three historical ETH/USD windows with annualised volatilities ranging from $41\%$ (calm, mid-2023) to $99\%$ (FTX crash, Nov 2022). [^9] USD used to sidestep fluctuations in stablecoin value which mean their volatility is not identically zero. The GBM-based $C(f)$ and $f^{*}$ track the realised costs well; in the volatile regime, heavier tails widen the confidence intervals but $f^{*}$ still sits near the observed minimum. The trending window has nonzero drift, which the driftless-GBM model does not capture. However, the retention ratio is price-independent (Proposition 1) regardless of drift, and the cross-term telescoping (Proposition 2) requires only that price increments are independent across blocks, not that they are driftless. Drift affects the LVR rate $\bar{\ell}$ but not the path optimality; over typical rebalancing windows (${\sim}8$ hours at $f^{*}$), drift is small relative to volatility.
8. Related Work
Dynamic weight AMMs
Several protocols implement G3M pools with time-varying weights. Balancer’s Liquidity Bootstrapping Pools [Martinelli and Mushegian, 2019] change weights on a fixed schedule to create a Dutch-auction-like price decline for token distribution; all deployed implementations (Balancer V1–V3, Fjord Foundry [Fjord Foundry, ]) use linear weight interpolation. Aera Finance V1 [Aera Finance (2022), Gauntlet] adjusts weights via external controllers, performing linear weight interpolation. In each case the weight schedule shapes the sequence of arbitrage opportunities the pool offers: an LBP is a Dutch auction designed to maximise price discovery, while a TFMM rebalancing is an auction on the total surplus paid to arbitrageurs. The interpolation curve governs the fine properties of the auction in both settings, but only the TFMM objective (cost minimisation) has been studied formally [Willetts and Harrington, 2024a].
Arbitrage cost and LVR
The LVR framework of Milionis et al. [2022] characterises the adverse-selection cost that arbitrageurs extract from AMM liquidity providers, with extensions to fee-adjusted settings [Milionis et al., 2023]. Evans [2019] analyses LP returns in G3Ms with time-varying weights. These results quantify the per-block cost of holding a position but do not optimise the schedule of weight changes within a rebalancing episode. TWAMM [White et al., 2021] solves the analogous problem for trade execution: a large swap is decomposed into infinitesimal sub-orders to reduce price impact, motivated by the same convexity that drives weight interpolation here.
Information geometry and portfolio theory
Pal and Wong [2017] study a geometry on the simplex induced by an L-divergence, with a generalised Pythagorean theorem governing optimal rebalancing frequency in stochastic portfolio theory. Their framework and ours both arrive at Riemannian structure on the weight simplex, but from different divergences for different reasons: L-divergence from portfolio excess growth, KL divergence from arbitrage cost.
9. Discussion
@@
In the language of information geometry [Amari, 1985, Amari and Nagaoka, 2000], the arbitrage cost is the canonical divergence of the dually-flat simplex (Appendix F). The Fisher–Rao metric provides a single framework that explains why the various candidate interpolation methods (linear, geometric, (AM+GM)/normalise, Lambert W) all perform similarly. Each corresponds to a different geodesic or approximate geodesic on the weight simplex (Appendix F); they nearly agree because the simplex is nearly flat at the scale of typical per-step weight changes. The sub-optimality bound (Theorem 9) makes this precise: for any reasonable step count, the choice between methods is practical, not a matter of loss optimality.
Accuracy of the quadratic approximation
The quadratic loss kernel, Eq (10), has cubic remainder in the per-step weight changes: for $f$-step interpolation the cumulative error is $O(1/f^{2})$. A $[1,1]$ Padé approximant resums this remainder into a rational function that matches the exact KL cost through $O(u^{4})$ and captures the cubic asymmetry between weight increases and decreases (Appendix M). The Lambert W approach of [Willetts and Harrington, 2024a] optimises each weight component independently and renormalises. It maximises the exact finite-step retention ratio rather than a quadratic surrogate, which makes it more accurate near the simplex boundary where the quadratic approximation is least tight (Appendix I).
Stochastic prices and practical deployment
The price-independence of the retention ratio (Proposition 1) means that a TFMM deployer can choose the interpolation curve without forecasting prices. The only new parameter is $f^{*}$, which requires an estimate of average LVR along the path; for a pool with known asset volatilities, this reduces to evaluating a single integral over the SLERP curve (Remark 1). On chains where fee income exceeds LVR extraction ($\phi<0$), even $f^{*}$ is unnecessary: the constant-price analysis applies and more steps is always preferable. If the target weights change mid-interpolation (e.g. a new strategy signal arrives before the current interpolation completes), the pool can restart SLERP from its current weights to the new target; since any sub-arc of a geodesic is itself a geodesic, the already-completed portion of the interpolation was optimal for the distance covered.
Fees and market microstructure
Swap fees and gas costs qualitatively change the tradeoff (§6). Fee revenue from arb trades telescopes along monotonic paths (Proposition 1), creating a path-independent offset that does not change the optimal interpolation curve (Corollary 2). The fee band caps the useful step count at $f_{\mathrm{threshold}}$ (Eq (27)), and the fee-adjusted optimum (Proposition 3) incorporates the empirical cost-of-carry $\phi$. On L2, where low gas costs enable high-frequency multi-hop routing at standard fees (e.g. 30 bp), $\phi$ can be negative, eliminating the finite-$f$ tradeoff entirely. The per-block weight change caps of [Willetts and Harrington, 2024c] bound the arbitrage opportunity available to an attacker with last-look/first-look across consecutive blocks; larger $f$ for the same total weight change means smaller per-block changes, which makes this attack harder.
@@
Computing SLERP requires one $\arccos$ evaluation, $N$ square roots, and per-step $\sin$ calls, comparable in cost to geometric mean interpolation. On-chain, where transcendental functions are expensive, (AM+GM)/normalise [Willetts and Harrington, 2024a] is still the pragmatic default. The recursive bisection algorithm (Corollary 7) shows that this heuristic computes exact geodesic points at power-of-two step counts using only elementary arithmetic. No trigonometric functions are needed.
10. Concluding remarks
The arbitrage cost of rebalancing a G3M pool is governed by the Fisher–Rao metric, and the optimal interpolation is its geodesic (SLERP), computable on-chain without trigonometric functions via recursive bisection. This structure survives stochastic prices, swap fees, and gas costs.
The geometric framework generalises beyond G3M pools. An AMM whose trading function is parameterised by a vector will typically have its arbitrage cost induce a Riemannian metric on its parameter space, and the optimal rebalancing trajectory will be a geodesic of that metric. For G3M, the metric is Fisher–Rao because the cost is a KL divergence; other designs (e.g. virtual balance interpolation in concentrated liquidity designs) would have different metrics and different geodesics. Identifying these metrics and their geodesics could yield analogous rebalancing strategies for other designs.
@@ Appendix A Retention Ratio
A.1. Deferred Proof: Proposition 1 [Retention ratio]
@@
(A.1) $R_{i}^{\mathrm{end}}=R_{i}^{\mathrm{start}}\frac{w_{i}^{\mathrm{end}}}{w_{i}^{\mathrm{start}}}\prod_{j=1}^{N}\left(\frac{w_{j}^{\mathrm{start}}}{w_{j}^{\mathrm{end}}}\right)^{w_{j}^{\mathrm{end}}},$
@@
(A.2) $\frac{R_{i}^{\mathrm{end}}}{R_{i}^{\mathrm{start}}}\frac{w_{i}^{\mathrm{start}}}{w_{i}^{\mathrm{end}}}=\prod_{j=1}^{N}\left(\frac{w_{j}^{\mathrm{start}}}{w_{j}^{\mathrm{end}}}\right)^{w_{j}^{\mathrm{end}}},$
and we thus define $r:=\prod_{j=1}^{N}\left(\frac{w_{j}^{\mathrm{start}}}{w_{j}^{\mathrm{end}}}\right)^{w_{j}^{\mathrm{end}}}$. For given reserves and market prices $\boldsymbol{\mathbf{m}}_{p}$, pool value is $V=\sum_{i=1}^{N}R_{i}m_{p,i}$. By the action of arbitrageurs, the pool holds the minimum value possible under its trading function. The weight vector then gives the division of value between assets [team, 2023a]: the G3M marginal price of token $i$ is $w_{i}k/R_{i}$ [Evans, 2019], which at equilibrium equals $p_{i}$, giving $R_{i}p_{i}=w_{i}V$ for all $i$. So we have
(A.3) $V^{\mathrm{start}}w^{\mathrm{start}}_{i}$ $=R^{\mathrm{start}}_{i}m_{p,i}$
(A.4) $V^{\mathrm{end}}w^{\mathrm{end}}_{i}$ $=R^{\mathrm{end}}_{i}m_{p,i}$
@@
(A.5) $\frac{V^{\mathrm{end}}}{V^{\mathrm{start}}}\frac{w^{\mathrm{end}}_{i}}{w^{\mathrm{start}}_{i}}$ $=\frac{R^{\mathrm{end}}_{i}}{R^{\mathrm{start}}_{i}}$
(A.6) $\Rightarrow\frac{V^{\mathrm{end}}}{V^{\mathrm{start}}}$ $=\frac{R_{i}^{\mathrm{end}}}{R_{i}^{\mathrm{start}}}\frac{w_{i}^{\mathrm{start}}}{w_{i}^{\mathrm{end}}}$
(A.7) $\Rightarrow\frac{V^{\mathrm{end}}}{V^{\mathrm{start}}}$ $=\prod_{j=1}^{N}\left(\frac{w_{j}^{\mathrm{start}}}{w_{j}^{\mathrm{end}}}\right)^{w_{j}^{\mathrm{end}}}$
(A.8) $\Rightarrow\frac{V^{\mathrm{end}}}{V^{\mathrm{start}}}$ $=r$
@@
A.2. Deferred Proof: Corollary 3 [Quadratic loss kernel]
See 3